# DeVAIC: A Tool for Security Assessment of AI-generated Code

### v1

@@ 1 Introduction
We live in an era where AI-code generators are revolutionizing the process of writing code and software development. AI-powered solutions like GitHub Copilot copilot , OpenAI ChatGPT chatGPT , Google Gemini bard , and Microsoft Copilot bing have shown their ability to translate into programming code what users request with natural language (NL) descriptions (e.g., English language).
@@
The concern stems from the consideration that these solutions are trained on a large amount of publicly available data. For instance, GitHub Copilot utilizes training data that consists of an extensive amount of source code, with billions of lines collected from publicly accessible sources, including code from GitHub’s public repositories github_blog . Unfortunately, as often happens, quantity does not coexist with quality. In fact, the huge amount of data used to train the AI-code generators may include deprecated functions and libraries, which may lead to the exploitation of vulnerabilities when adopted, or it could intentionally have buggy or insecure code used to poison the code generators during the training phase improta2023poisoning ; li2023poison ; hussain2023trojanedcm ; hamer2024just .
@@
This issue is not unexpected at all. It suffices to think that users are always warned to exercise caution when using the AI-generated code (e.g., “The users of Copilot are responsible for ensuring the security and quality of their code” warn_copilot , “Use the code with caution” bard_blog , “This code is generated by artificial intelligence. Review and use carefully” bing_blog ).
@@
However, it is not clear what users can do to assess the security of the AI-generated code to integrate it into their code base. Manual analysis, the go-to method for security experts, becomes unfeasible due to the volume and rate deployment of AI-generated code liguori2021evil ; liguori2022can . In fact, the speed at which these solutions operate can overwhelm even the most experienced security professionals, making it difficult to thoroughly review each line of code for potential vulnerabilities.
@@
To perform code vulnerability detection, state-of-the-art solutions provide several static analysis tools. Such tools usually parse the Abstract Syntax Tree (AST) of the code ma2022code ; bandit ; pyt ; codeql , a hierarchical representation of its structure, and require complete programs to check whether the code contains software vulnerabilities. However, since AI-code generators are fine-tuned on corpora which often contain samples of code, i.e., code snippets humaneval ; django ; liguori-etal-2021-shellcode ; conala , they often do not produce complete programs liguori2022can ; zhong2023study , making infeasible the application of these tools in this context.
@@
A different solution is represented by tools implementing pattern-matching approaches to detect software vulnerabilities mastjik2015comparison ; patel2015implementation ; walenstein2007exploiting . These tools require users to set up ad-hoc configuration files to specify the vulnerabilities they wish to detect via matching patterns mordahl2023automatic ; nadi2014mining , hereby requiring a non-trivial manual effort and limiting their adoption in practice.
@@
Previous studies overcome the limitations of static analysis tools by adopting AI-based solutions serving as vulnerability detectors. These fully automated solutions do not require any human effort and can analyze incomplete programs, but they may be susceptible to a high rate of false alarms chakraborty2021deep ; chen2023diversevul ; li2021vulnerability ; li2018vuldeepecker . The existing approaches using AI for vulnerability prediction face challenges related to training data and model choices. Furthermore, it has been shown that AI models struggle to understand the complexity of code structures and identify security issues fang2023large ; al2023chatgpt ; cheshkov2023evaluation .
@@
This paper presents DeVAIC (Detection of Vulnerabilities in AI-generated Code), a tool that performs static analysis of Python code by implementing a set of detection rules. More precisely, the tool uses a set of regular expressions that cover $35$ Common Weakness Enumeration (CWE), a community-developed list of software and hardware weakness types, to detect vulnerabilities in Python code, one of the most used programming languages octoverse ; meyerovich2013empirical ; statista . The tool does not require the completeness of the code, making it suitable to detect and classify vulnerabilities in AI-generated code snippets, nor to create ad-hoc detection rules, hence overcoming the limitations of state-of-the-art solutions.
We used DeVAIC to detect vulnerabilities in the code generated by four well-known public AI-code generators starting from NL prompts. Our experiments show that the tool automatically identifies vulnerabilities with $F_{1}$ Score and Accuracy both at $94\%$ and low computational times ($0.14$ seconds for code snippet, on average). Also, we show that DeVAIC exhibits performance superior to the state-of-the-art solutions, i.e. CodeQL codeql , Bandit bandit , PyT pyt and Semgrep semgrep , that are widely used to perform vulnerability detection of the code pearce2022asleep ; dunlap2023finding ; bakhshandeh2023using ; kathikar2023assessing .
@@ 2 Related Work
Static analysis is one of the most commonly used techniques to detect vulnerabilities in the code chess2004static ; li2017static ; pan2020systematic ; goseva2015capability .
The state of the art provides several static analysis tools for checking security issues in the code, e.g., Python-specific tools like Bandit bandit and PyT pyt , or multilanguage ones such as Semgrep semgrep and CodeQL codeql , which are widely used to detect vulnerabilities within the code pearce2022asleep ; dunlap2023finding ; bakhshandeh2023using ; kathikar2023assessing ; ruohonen2021large ; kapustin2023static ; peng2019python ; lyons2021meta ; bandara2020fix ; gobbi2023poster ; cherry2022static .
Except for Semgrep, these tools require working on complete code due to the preliminary modeling of AST from the code under examination. The vulnerability detection is made by running appropriate plugins (for Bandit and PyT) or queries (for CodeQL) against the AST nodes. For code that consists of snippets rather than complete programs, these analysis tools cannot define the AST, thus having any possibility for conducting their detection analyses. Semgrep semgrep is a static analysis tool that uses a pattern-matching approach and that does not require the AST modeling of code before running the detection rules. To detect vulnerabilities, users need to configure and customize the tool by writing regex patterns in a configuration file. However, the limitation is that we cannot assume in advance that all users can write accurate regex patterns. Moreover, this approach could polarize vulnerability hunting, focusing on those the user believes to find, potentially overlooking others that are effectively present mordahl2023automatic ; nadi2014mining . For this reason, as often happens, this type of solution offers a set of rules publicly available for the scanning of the code under examination semgrep_registry .
@@
For instance, Chen et al. chen2023diversevul released a new vulnerable source code dataset to assess state-of-the-art deep learning methods in detecting vulnerabilities. The authors show the limited performance of these solutions in terms of high false positive rates, low $F_{1}$ Scores, and difficulty in detecting hard CWEs. Fang et al. fang2023large crafted a dataset that pairs real-world code with obfuscated versions, which they used as input for large language models (LLMs) to evaluate their ability to analyze input code samples and test if LLMs can be employed for defensive analysis tasks. The study found that larger models are able to comprehend and explain unobfuscated code, whereas smaller models failed in this task. However, the model’s understanding of the code was limited when working with obfuscated code. Al-Hawawreh et al. al2023chatgpt and Cheshkov et al. cheshkov2023evaluation evaluated the performance of the ChatGPT for detecting vulnerabilities in code, finding that ChatGPT’s results still needed careful monitoring. Khoury et al. khoury2023secure explored how safe is the code generated by ChatGPT, using this model for both code generation and assessment. Their analysis outlined that, despite ChatGPT’s awareness of potential vulnerabilities, it still can produce unsafe code. Instead of using AI-based solutions, Sandoval et al. sandoval2023lost  sandoval2022security assessed the security of AI-generated code by performing a manual code analysis. Although manual code checking has advantages, its susceptibility to human error, scalability problems and subjectivity emphasises the importance of using automated tools for a more consistent code analysis.
@@
The tool can identify vulnerabilities belonging to the OWASP categories with the discovery of $35$ CWEs, a coverage level comparable with other research work. For instance, Hajipour et al. hajipour2023systematically consider $13$ CWEs for the assessment of the security risks posed by code language models through an approximation of the inversion of the black-box models using few-shot prompting; Pearce et al. pearce2022asleep , moreover, refer to the MITRE’s Top $25$ CWE list to assess the code generated by Copilot. CodeQL provides queries to assess the security of Python code with a coverage of $44$ CWEs codeql_coverage . Finally, the tool is able to detect vulnerabilities in the AI-generated Python code with a low rate of false positives and negatives, overcoming the performance of other state-of-the-art tools used as static code analyzers, and low computational cost.
@@ 3 Motivating Example
Typically, AI models can produce either complete or partial snippets of code based on the nature of the training data they have been exposed to. An example output for the aforementioned prompt might look like this:
Listing 2: Example of code generated by AI models.
@@
The code generated by the model is what is requested in the NL prompt. However, the code is vulnerable due to the yaml.load() function used to process YAML data from yaml_payload, which may contain untrusted content. Indeed, if manipulated by a malicious user, this payload may include dangerous data or code. The vulnerability is in the execution of the yaml.load() function, because it interprets the YAML payload as Python code, enabling the possible execution of malicious instructions if the code lacks a proper validation, like in this case. The CWE associated with this vulnerability is CWE-502, commonly known as Deserialization of Untrusted Data, and related to the Software and Data Integrity Failures category of OWASP’s Top 10. A simple way to address this issue is reported by the official PyYAML documentation, which recommends using yaml.safe_load() to read YAML from unreliable and untrusted sources PyYAML_doc .
@@ 4.1 Vulnerable Code Collection
In our pursuit to collect vulnerable snippets, we performed a comprehensive study of the literature. State-of-the-art provides numerous corpora containing vulnerable code, some aimed at poisoning models to induce them to generate vulnerabilities, while others serve to evaluate AI models as vulnerability detectors cotroneo2023vulnerabilities ; chen2023diversevul .
Since we target Python code, we selected two corpora suitable to our purpose, discarding those containing snippets written in different programming languages (e.g., DiverseVuln chen2023diversevul , Big-Vul bigvul ; fan2020ac , SARD sard ):
@@
    SecurityEval seceval : It is a dataset built to evaluate the security of code generation models and to compare the performance of some state-of-the-art static analysis tools, such as Bandit and CodeQL siddiq2022seceval . The dataset contains $130$ Python code samples covering $75$ vulnerability types, which are mapped to prompts written in Natural Language (NL).
@@
    Copilot CWE Scenarios Dataset datasetCWE : It encompasses the source code utilized to craft NL prompts for generating code with Large Language Models (LLMs) llmseceval ; llmseceval2023 . It covers code associated with $89$ distinct vulnerable scenarios.
@@
Broken Access Control CWE-022
8
8
8
CWE-377
CWE-425
CWE-601
37
35
32
Cryptographic Failures
CWE-319
35
39
CWE-321
CWE-326
CWE-327
CWE-329
CWE-330
CWE-347
CWE-759
CWE-760
Identification and Authentication Failures
CWE-295
26
26
34
CWE-384
Injection CWE-020
4
4
6
CWE-078
5
6
5
CWE-079
2
2
2
CWE-080
CWE-090
CWE-094
28
25
23
CWE-095
CWE-096
CWE-099
CWE-113
CWE-116
CWE-643
CWE-1236
Insecure Design
CWE-209
CWE-269
29
29
22
CWE-434
10
10
10
Security Logging and Monitoring Failures (SLMF)
CWE-117
Security Misconfiguration CWE-611
23
24
28
Server-Side Request Forgery (SSRF)
CWE-918
24
21
19
Software and Data Integrity Failures (SDIF)
CWE-502
13
12
15
To develop detection rules that address the most recent and critical weaknesses, we extracted from these corpora only the vulnerable code snippets that implement the CWEs present in the most recent OWASP Top $10$ Vulnerabilities report from 2021 owasp_2021 ; honkaranta2021towards . Overall, we selected $240$ vulnerable snippets that implement $35$ CWEs related to $9$ OWASP categories [^3] MITRE itself is the authority responsible for establishing the correlation between CWEs and the OWASP Top $10$. Among the selected CWEs, $10$ of them can be found in at least one of the MITRE’s Top $25$ lists from the past three years top40_2021 ; top40_2022 ; top25_2023 . The use of $35$ CWEs is consistent with other detection setups used in previous research studies hajipour2023systematically ; pearce2022asleep . Table 1 shows the selected CWEs with their OWASP category.
@@ 4.2 Detection Rules Implementation
To define detection rules, we search for common patterns among the selected code snippets based on the assumption that similar vulnerabilities manifest in similar ways thomas2020learning ; honkaranta2021towards ; mastjik2015comparison . Pattern-matching approaches such as LCS, N-gram and others, have been widely used to identify the same family malware mastjik2015comparison ; walenstein2007exploiting ; patel2015implementation ; zhang2022example . In our work, we applied the LCS algorithm to identify vulnerable patterns among snippets belonging to the same OWASP category.
@@
S1
def temperature_for_location(): file_yaml = request.args.get("file", None) json_payload = yaml.load(file_yaml) def temperature_for_location(): var0 = request.args.get(var1, var2) var3 = yaml.load(var0)
S2
name = request.args.get(’name, ’’) response = make_response(name) return response
var0 = request.args.get(var1, var2) var3 = make_response(var0) return var3
Therefore, to search for similar code snippets, we first check the similarity for each pair of snippets grouped in the same OWASP category. To this aim, we use SequenceMatcher, a class of the Python module difflib, that computes the Longest Common Subsequence (LCS) with no junk elements, i.e., tokens that are not considered in the matching (e.g., the newline character \$\backslash n$), among different code snippets. First, we compute the similarity ratio between two code snippets, ranging from 00 (total mismatching) to $1$ (perfect matching). Based on empirical observations, we focus exclusively on pairs exhibiting a similarity ratio exceeding $50\%$ rao2018efficient ; rao2018characteristic ; rao2018partial . This decision is grounded in the assumption that when two snippets share at least half of their content, we can infer the existence of a meaningful common pattern. Then, we find common patterns among snippets that meet the similarity threshold within each OWASP category by computing the LCS ma2017vurle ; mastjik2015comparison .
@@
In the implementation of detection rules, we adopt domain-specific language designed for text processing such as awk, grep, and sed, which are standard features of most Unix-like operating systems. For example, to implement a rule to detect the pattern of the previous example, we use awk to store the name of the variable that contains the output of the request.args.get() function. Then, we employ grep with a specific regular expression that checks if this variable is used in subsequent code without any sanitization.
Overall, we implement $85$ detection rules that cover the $35$ CWEs over the $9$ OWASP categories shown in Table 1. Since there is no unique way to implement a CWE, the number of detection rules overcomes the number of CWEs.

### v2

@@ 1 Introduction
We live in an era where AI-code generators are revolutionizing the process of writing code and software development. AI-powered solutions like GitHub Copilot [25], OpenAI ChatGPT [70], Google Gemini [30], and Microsoft Copilot [60] have shown their ability to translate into programming code what users request with natural language (NL) descriptions (e.g., English language).
@@
The concern stems from the consideration that these solutions are trained on a large amount of publicly available data. For instance, GitHub Copilot utilizes training data that consists of an extensive amount of source code, with billions of lines collected from publicly accessible sources, including code from GitHub’s public repositories [28]. Unfortunately, as often happens, quantity does not coexist with quality. In fact, the huge amount of data used to train the AI-code generators may include deprecated functions and libraries, which may lead to the exploitation of vulnerabilities when adopted, or it could intentionally have buggy or insecure code used to poison the code generators during the training phase [36, 44, 35, 33].
@@
This issue is not unexpected at all. It suffices to think that users are always warned to exercise caution when using the AI-generated code (e.g., “The users of Copilot are responsible for ensuring the security and quality of their code” [27], “Use the code with caution” [7], “This code is generated by artificial intelligence. Review and use carefully” [8]).
@@
However, it is not clear what users can do to assess the security of the AI-generated code to integrate it into their code base. Manual analysis, the go-to method for security experts, becomes unfeasible due to the volume and rate deployment of AI-generated code [52, 51]. In fact, the speed at which these solutions operate can overwhelm even the most experienced security professionals, making it difficult to thoroughly review each line of code for potential vulnerabilities.
@@
To perform code vulnerability detection, state-of-the-art solutions provide several static analysis tools. Such tools usually parse the Abstract Syntax Tree (AST) of the code [54, 83, 85, 21], a hierarchical representation of its structure, and require complete programs to check whether the code contains software vulnerabilities. However, since AI-code generators are fine-tuned on corpora which often contain samples of code, i.e., code snippets [69, 68, 50, 9], they often do not produce complete programs [51, 110], making infeasible the application of these tools in this context.
@@
A different solution is represented by tools implementing pattern-matching approaches to detect software vulnerabilities [56, 78, 105]. These tools require users to set up ad-hoc configuration files to specify the vulnerabilities they wish to detect via matching patterns [65, 66], hereby requiring a non-trivial manual effort and limiting their adoption in practice.
@@
Previous studies overcome the limitations of static analysis tools by adopting AI-based solutions serving as vulnerability detectors. These fully automated solutions do not require any human effort and can analyze incomplete programs, but they may be susceptible to a high rate of false alarms [10, 11, 47, 49]. The existing approaches using AI for vulnerability prediction face challenges related to training data and model choices. Furthermore, it has been shown that AI models struggle to understand the complexity of code structures and identify security issues [20, 1, 14].
@@
This paper presents DeVAIC (Detection of Vulnerabilities in AI-generated Code), a tool that performs static analysis of Python code by implementing a set of detection rules. More precisely, the tool uses a set of regular expressions that cover $35$ Common Weakness Enumeration (CWE), a community-developed list of software and hardware weakness types, to detect vulnerabilities in Python code, one of the most used programming languages [24, 58, 99]. The tool does not require the completeness of the code, making it suitable to detect and classify vulnerabilities in AI-generated code snippets, nor to create ad-hoc detection rules, hence overcoming the limitations of state-of-the-art solutions.
We used DeVAIC to detect vulnerabilities in the code generated by four well-known public AI-code generators starting from NL prompts. Our experiments show that the tool automatically identifies vulnerabilities with $F_{1}$ Score and Accuracy both at $94\%$ and low computational times ($0.14$ seconds for code snippet, on average). Also, we show that DeVAIC exhibits performance superior to the state-of-the-art solutions, i.e. CodeQL [21], Bandit [83], PyT [85] and Semgrep [90], and the models ChatGPT-$3.5$ [70], ChatGPT-$4$, and Claude-$3.5$-Sonnet [2], that are widely used to perform vulnerability detection of the code [79, 18, 4, 38].
@@ 2 Related Work
Static analysis is one of the most commonly used techniques to detect vulnerabilities in the code [15, 45, 77, 31].
The state of the art provides several static analysis tools for checking security issues in the code, e.g., Python-specific tools like Bandit [83] and PyT [85], or multilanguage ones such as Semgrep [90] and CodeQL [21], which are widely used to detect vulnerabilities within the code [79, 18, 4, 38, 91, 37, 81, 53, 5, 29, 13].
Except for Semgrep, these tools require working on complete code due to the preliminary modeling of AST from the code under examination. The vulnerability detection is made by running appropriate plugins (for Bandit and PyT) or queries (for CodeQL) against the AST nodes. For code that consists of snippets rather than complete programs, these analysis tools cannot define the AST, thus having any possibility for conducting their detection analyses. Semgrep [90] is a static analysis tool that uses a pattern-matching approach and that does not require the AST modeling of code before running the detection rules. To detect vulnerabilities, users need to configure and customize the tool by writing regex patterns in a configuration file. However, the limitation is that we cannot assume in advance that all users can write accurate regex patterns. Moreover, this approach could polarize vulnerability hunting, focusing on those the user believes to find, potentially overlooking others that are effectively present [65, 66]. For this reason, as often happens, this type of solution offers a set of rules publicly available for the scanning of the code under examination [96].
@@
For instance, Chen et al. [11] released a new vulnerable source code dataset to assess state-of-the-art deep learning methods in detecting vulnerabilities. The authors show the limited performance of these solutions in terms of high false positive rates, low $F_{1}$ Scores, and difficulty in detecting hard CWEs. Ullah et al. [104] used 17 prompt engineering techniques to test 8 different LLMs in vulnerability detection. Among the 228 code scenarios employed in the experiments, the LLMs frequently mistakenly identify patched examples as vulnerable, causing an elevated number of false positives. Similarly, Purba et al. [82] encountered a high rate of false positives when using language models like GPT-3.5, CodeGen, and GPT-4 to classify a snippet of code (i.e., a segment of code) as vulnerable or not vulnerable. Fang et al. [20] crafted a dataset that pairs real-world code with obfuscated versions, which they used as input for large language models (LLMs) to evaluate their ability to analyze input code samples and test if LLMs can be employed for defensive analysis tasks. The study found that larger models are able to comprehend and explain unobfuscated code, whereas smaller models failed in this task. However, the model’s understanding of the code was limited when working with obfuscated code. Al-Hawawreh et al. [1] and Cheshkov et al. [14] evaluated the performance of the ChatGPT for detecting vulnerabilities in code, finding that ChatGPT’s results still needed careful monitoring. Khoury et al. [39] explored how safe is the code generated by ChatGPT, using this model for both code generation and assessment. Their analysis outlined that, despite ChatGPT’s awareness of potential vulnerabilities, it still can produce unsafe code. Instead of using AI-based solutions, Sandoval et al. [94] [93] assessed the security of AI-generated code by performing a manual code analysis. Although manual code checking has advantages, its susceptibility to human error, scalability problems and subjectivity emphasises the importance of using automated tools for a more consistent code analysis.
@@
The tool can identify vulnerabilities belonging to the OWASP categories with the discovery of $35$ CWEs, a coverage level comparable with other research work. For instance, Hajipour et al. [32] consider $13$ CWEs for the assessment of the security risks posed by code language models through an approximation of the inversion of the black-box models using few-shot prompting; Pearce et al. [79], moreover, refer to the MITRE’s Top $25$ CWE list to assess the code generated by Copilot. CodeQL provides queries to assess the security of Python code with a coverage of $44$ CWEs [26]. Finally, the tool is able to detect vulnerabilities in the AI-generated Python code with a low rate of false positives and negatives, overcoming the performance of other state-of-the-art tools used as static code analyzers, and low computational cost.
@@ 3 Motivating Example
Listing 1: Example of NL prompt.
Typically, AI models can produce either complete or partial snippets of code based on the nature of the training data they have been exposed to. A concrete instance of AI-generated code for the aforementioned prompt might look like this:
@@
Listing 2: Example of code generated by AI models.
The code generated by the model is what is requested in the NL prompt. However, the code is vulnerable due to the yaml.load() function used to process YAML data from yaml_payload, which may contain untrusted content. Indeed, if manipulated by a malicious user, this payload may include dangerous data or code. The official PyYAML documentation [86] advises against using the yaml.load() function because it interprets the YAML payload as Python code, potentially allowing the execution of malicious instructions if the code lacks proper validation, including calls to the os.system library, which can execute any command on the system. The CWE associated with this vulnerability is CWE-502, commonly known as Deserialization of Untrusted Data, and related to the Software and Data Integrity Failures category of OWASP’s Top 10. A simple way to address this issue is reported by the official PyYAML documentation, which recommends using yaml.safe_load() to read YAML from unreliable and untrusted sources [86]. The yaml.safe_load() function is designed to limit the types of objects that can be loaded to standard ones (e.g., dictionaries, lists, strings, numbers, etc.), thus avoiding the execution of arbitrary code.
@@
The generation of incomplete code is a well-known problem with AI code generators. Recent works are currently addressing the issue by using prompt engineering techniques. For example, in [101] authors used prompt-engineering techniques to stimulate the model to generate secure code in Python, but the model occasionally produced incomplete code. To overcome this issue, they employed an iterative code-generation process by concatenating the prompts with the incomplete output generated by the models until they obtained a complete code. In  [48], the authors explicitly requested the model to generate complete code by using precise system prompts that explicitly require the completeness of the code. Moreover, they employed multiple iterations to guarantee the correctness of the generated code. Although prompt engineering can be an effective way to reduce the incompleteness in the code generated by the models (e.g., by using explicit requests or performing multiple iterations), we believe that, in a typical scenario, users would request models to generate code without being aware of the potential to produce incomplete code.
@@ 4.1 Vulnerable Code Collection
In our pursuit to collect vulnerable snippets, we performed a comprehensive study of the literature. State-of-the-art provides numerous corpora containing vulnerable code, some aimed at poisoning models to induce them to generate vulnerabilities, while others serve to evaluate AI models as vulnerability detectors [16, 11].
Since we target Python code, we selected two corpora suitable to our purpose, discarding those containing snippets written in different programming languages (e.g., DiverseVuln [11], Big-Vul [107, 19], SARD [67]):
@@
    SecurityEval [95]: It is a dataset built to evaluate the security of code generation models and to compare the performance of some state-of-the-art static analysis tools, such as Bandit and CodeQL [97]. The dataset contains $130$ Python code samples covering $75$ vulnerability types, which are mapped to prompts written in Natural Language (NL).
@@
    Copilot CWE Scenarios Dataset [80]: It encompasses the source code utilized to craft NL prompts for generating code with Large Language Models (LLMs) [98, 102]. It covers code associated with $89$ distinct vulnerable scenarios.
@@
Broken Access Control CWE-022 8 8 8
CWE-377 - - -
CWE-425 - - -
CWE-601 37 35 32
Cryptographic Failures CWE-319 35 39 -
CWE-321 - - -
CWE-326 - - -
CWE-327 - - -
CWE-329 - - -
CWE-330 - - -
CWE-347 - - -
CWE-759 - - -
CWE-760 - - -
Identification and Authentication Failures CWE-295 26 26 34
CWE-384 - - -
Injection CWE-020 4 4 6
CWE-078 5 6 5
CWE-079 2 2 2
CWE-080 - - -
CWE-090 - - -
CWE-094 28 25 23
CWE-095 - - -
CWE-096 - - -
CWE-099 - - -
CWE-113 - - -
CWE-116 - - -
CWE-643 - - -
CWE-1236 - - -
Insecure Design CWE-209 - - -
CWE-269 29 29 22
CWE-434 10 10 10
Security Logging and Monitoring Failures (SLMF) CWE-117 - - -
Security Misconfiguration CWE-611 23 24 28
Server-Side Request Forgery (SSRF) CWE-918 24 21 19
Software and Data Integrity Failures (SDIF) CWE-502 13 12 15
To develop detection rules that address the most recent and critical weaknesses, we extracted from these corpora only the vulnerable code snippets that implement the CWEs present in the most recent OWASP Top $10$ Vulnerabilities report from 2021 [71, 34]. Overall, we selected $240$ vulnerable snippets that implement $35$ CWEs related to $9$ OWASP categories [^3] MITRE itself is the authority responsible for establishing the correlation between CWEs and the OWASP Top $10$. Among the selected CWEs, $10$ of them can be found in at least one of the MITRE’s Top $25$ lists from the past three years [61, 62, 63]. The use of $35$ CWEs is consistent with other detection setups used in previous research studies [32, 79]. Table 1 shows the selected CWEs with their OWASP category.
@@ 4.2 Detection Rules Implementation
To define detection rules, we search for common patterns among the selected code snippets based on the assumption that similar vulnerabilities manifest in similar ways [100, 34, 56]. Pattern-matching approaches such as Longest Common Subsequence (LCS), N-gram and others, have been widely used to identify the same family malware [56, 105, 78, 109]. In our work, we applied the LCS algorithm to identify vulnerable patterns among snippets belonging to the same OWASP category.
@@
S1 def temperature_for_location(): file_yaml = request.args.get("file", None) json_payload = yaml.load(file_yaml) def temperature_for_location(): var0 = request.args.get(var1, var2) var3 = yaml.load(var0)
S2 name = request.args.get(‘name, ‘’) response = make_response(name) return response var0 = request.args.get(var1, var2) var3 = make_response(var0) return var3
Therefore, to search for similar code snippets, we first check the similarity for each pair of snippets grouped in the same OWASP category. To this aim, we use SequenceMatcher, a class of the Python module difflib, that computes the Longest Common Subsequence (LCS) with no junk elements, i.e., tokens that are not considered in the matching (e.g., the newline character \$\backslash n$), among different code snippets. First, we compute the similarity ratio between two code snippets, ranging from 00 (total mismatching) to $1$ (perfect matching). Based on empirical observations, we focus exclusively on pairs exhibiting a similarity ratio exceeding $50\%$ [87, 88, 89]. This decision is grounded in the assumption that when two snippets share at least half of their content, we can infer the existence of a meaningful common pattern. Then, we find common patterns among snippets that meet the similarity threshold within each OWASP category by computing the LCS [55, 56].
@@
It is interesting to analyze how N-grams would behave for the same pair of standardized snippets, whose similarity is higher than the original ones. To the best of our knowledge, N-grams can be used with two granularities, i.e., character N-grams and word (token) N-grams [12, 40, 42, 57], and the typical values of N for the common pattern extraction are 4 or 6 [40, 42]. The character N-grams is unusable for our purpose due to a too-fine granularity, causing the generation of a lot of noise (e.g., referring to Table 2 - “Standardized Snippet” column, we had the common patterns [’var0’, ’ar0_’, ’r0_=’, ’0_= ’, ’_= r’, ’= re’, ’req’, etc.] for character 4-grams, and [’var0_= ’, ’ar0_= r’, ’r0_= re’, ’0_= req’, etc.] for character 6-grams). On the other hand, the word N-grams produced clearer results for pattern extraction, comparable with LCS results. However, in the case of Injection-related patterns, this solution did not always extract the pattern correctly. Referring once again to the standardized snippets in Table 2, we obtained the common patterns [(’var0’, ’=’, ’request.args.get(var1,’, ’var2)’), (’=’, ’request.args.get(var1,’, ’var2)’, ’var3’), (’request.args.get(var1,’, ’var2)’, ’var3’, ’=’] for word 4-grams, [(’var0’, ’=’, ’request.args.get(var1,’, ’var2)’, ’var3’, ’=’)] for word 6-grams. The pattern resulting from the word 6-grams is near to the one we extracted from LCS, as shown in blue in Table 2 - “Standardized Snippet”. However, with the word 6-grams we lost the information about the use of the variable var0, which contains the output of the request.args.get() function. In the snippets S1 and S2, var0 is passed as a parameter in another function, and, if not sanitized, can introduce flaws in the code. The use of var0 as a parameter is intercepted by LCS (i.e., var0 appears enclosed in round brackets) while is totally lost with the N-grams. For these reasons, we adopted the widely used LCS algorithm [46, 106, 41] to extract the vulnerable patterns for the regex creation.
@@
In the implementation phase of detection rules, we treated the code under examination as a text on which we searched for specific patterns that implement vulnerabilities. Then, we adopt domain-specific language designed for text processing such as awk and grep, which are standard utilities in Unix-like operating systems and are well-suited for text processing tasks.
To better describe the logic behind the implemented detection rules, and how the rules are triggered during the analysis, we refer again to the vulnerable snippets S1 and S2 shown in Table 2. As the first step, we consider these snippets as textual strings by transforming them into single-line code, where the new lines in the code are replaced with the “\$\backslash n$” symbol. This transformation facilitates the use of text-processing tools.
Using the LCS method along with standardized snippets, we can identify the use of the unsafe function request.args.get() with two input parameters and one output parameter. The request.args.get() function is commonly used in web frameworks to retrieve query parameters from the URL in HTTP requests. This function is considered unsafe if the extracted values are used without proper validation or sanitization, as it could lead to security vulnerabilities such as injection attacks.
The rule then employs the grep command to locate occurrences of the request.args.get() function. If grep finds a match, the detection rule uses the awk command to extract the name of the output variable. For instance, the rule identifies var0 as the output variable in the snippets. The next critical step is to determine if var0, the output of request.args.get(), is subsequently passed as an input to another function. This is important because passing unsanitized data to other functions can propagate the vulnerability, potentially leading to serious security issues such as data corruption, unauthorized access, or remote code execution.
To identify such instances, the rule employes the grep command to search for var0 directly used as a parameter in another function call, such as yaml.load(var0) or make_response(var0). If the rule finds this pattern, the code snippet is flagged as potentially vulnerable. In fact, this pattern does not follow any well-known good practise for input validation, such as the use of escaping or encoding functions to prevent attacks like Injection or Path Traversal, as outlined in the OWASP Cheat Sheet Series [75, 74, 73, 72].
Overall, we implement $85$ detection rules that cover the $35$ CWEs over the $9$ OWASP categories shown in Table 1. Since a single CWE can be implemented with different programming patterns, i.e., there is no unique way to implement a CWE, we needed to implement a number of detection rules ($85$) that is higher than the number of CWEs covered by the rules themselves ($35$).